# A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability

### v2

@@ 4.1.4. Trust Assumptions
The solutions that introduce trust assumptions do it for different purposes. The approaches using zk-SNARK as ZKP require a trusted setup to initialize and correctly share some cryptographic parameters (e.g., (Miers et al., 2013; Garman et al., 2017; Ben Sasson et al., 2014; Xue et al., 2023)). A few works require a trusted central bank to issue coins that can be later exchanged anonymously (e.g., (Goodell et al., 2023; Friolo et al., 2024)). Elfadul et al. (Elfadul et al., 2024) assume the presence of a regulatory supervision that grants permissions to users to perform transaction; this centralized authority could asymmetrically revoke such authorization. Contributions that explore special forms of auditability related to revocable anonymity (e.g. (Wüst et al., 2019; Lin et al., 2020; Xue et al., 2023; Yu et al., 2024)), trapdoor (e.g., (Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020; Elfadul et al., 2024)), or user tracing (e.g., (Garman et al., 2017)) require trust in the tracing authority that uncovers transaction details only when necessary. PEReDi (Kiayias et al., 2022) reduces trust requirements by relying on quorum of centralized entities that must agree to perform specific operations (such as decrypting transactions).
@@ 4.1.5. Interaction Models
Privacy-preserving digital payment systems differ in whether transactions require direct interaction between payer and payee. Interactive payment protocols need active cooperation during the transaction process. This is common in account-based models where recipients help in constructing or approving the updated account state, via ZKPs or digital signatures. For example, PEReDi (Kiayias et al., 2022) and Platypus (Wüst et al., 2022) enforce synchronized interaction: both parties participate in updating the ledger with proofs and commitments, ensuring mutual accountability and preventing attacks like dusting or replay. Similarly, Chaumian token systems often require immediate token claiming to prevent double spending (e.g., (Dold, 2019)). In contrast, non-interactive payment protocols let the payer unilaterally create and broadcast transactions to the network. Many token-based systems, especially unspent token models, follow this approach (e.g., (Ben Sasson et al., 2014; Van Saberhagen, 2013; Noether et al., 2016; Poelstra, 2016)). This improves user convenience and enables asynchronous payments, although receivers may need periodic ledger scanning or local computation. [^8] Off-chain communication can alleviate the burden of scanning the ledger, but current protocols lack this feature. Some account-based protocols approximate non-interactive behavior via delayed claiming. For instance, Sarencheh et al. (Sarencheh et al., 2025) propose a two stage and non-interactive variant of PEReDi: first, the sender interacts with issuers to update his account (transfer stage); then, the receiver claims funds using enclosed ZKP and credits his account (claim stage). Overall, interactive designs enable rich policy enforcement and conditional privacy, whereas non-interactive models better suit high-latency or asynchronous networks.
@@ 4.2.1. Ledger Data Models
Account-based ledgers associate a balance with each account address, which is debited or credited by transactions. This simplifies state management and enables simple definitions of policies based on cumulative balances. Nonetheless, it requires safeguards against replay and double-spending attacks, often implemented by storing a sequence number (or nonce) in each account and updating it with every transaction. Since the balance of an account is explicitly represented, obtaining anonymity and unlinkability is challenging, especially in systems where users cannot freely generate multiple addresses (e.g., (Guan et al., 2022)). Privacy-preserving account-based solutions replace plain-text balances with commitments, achieving confidentiality of balances and transactions (e.g., (Fauzi et al., 2019; Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022; Sarencheh et al., 2025)). Liu et al. (Liu et al., 2022) propose an account-based CBDC with sender and receiver anonymity but with no unlinkability. The system follows a two-tier model where the central bank can identify users via encrypted identifiers, while commercial banks cannot. To guarantee anonymity while ensuring unlinkability, most of the approaches change users’ account address at each transaction and use an authenticator to prove correctness. The authenticator can be implemented, e.g., using digital signatures by trusted or centralized authorities (e.g., (Kiayias et al., 2022)) or ZKPs (e.g., (Wüst et al., 2022; Fauzi et al., 2019; Bünz et al., 2020; Papadoulis et al., 2024)). Platypus (Wüst et al., 2022) rely on ZKPs; conversely, in PEReDi (Kiayias et al., 2022), account states must be (blindly) signed by a quorum of maintainers, and since each transaction reveals prior account snapshots to construct new authenticators, randomized signatures are used to prevent linkability. Quisquis (Fauzi et al., 2019) and AQQUA (Papadoulis et al., 2024) use updatable public keys to update accounts in a decentralized manner. The account-based model lends itself well to auditability, as it allows simple definition of policies on holding limits (e.g., (Gross et al., 2021; Wüst et al., 2022; Papadoulis et al., 2024)) or spending and receiving limits (e.g., (Papadoulis et al., 2024)). Gross et al. (Gross et al., 2021) includes, in the account state, an accumulator of all spending in the current epoch, enabling policies on this information. AQQUA (Papadoulis et al., 2024) includes commitments to three different counters (i.e., balance, total amounts sent and received). Conversely, Platypus (Wüst et al., 2022) leaves the specification of the additional information to be application dependent.
@@
Chaumian tokens (Chaum, 1983) use blind signatures and were first deployed in centralized, bank-issued electronic cash schemes. When issuing a token, the bank blindly signs a random value provided by the payer, allowing the payer to later unblind it for spending (ensuring sender anonymity). These tokens rely on a trusted authority to issue and verify their validity; to this end, the authority must hold an ever increasing list of spent tokens that might introduce scalability challenges. Some implementations introduce expiration to limit circulation (e.g., (Dold, 2019; Lovejoy et al., 2023)). Chaumian-inspired decentralized privacy-preserving systems (e.g., Monero (Noether et al., 2016), Zerocoin (Miers et al., 2013), and Zerocash (Ben Sasson et al., 2014)) eliminate the need of a central issuer using a combination of different cryptographic techniques. For example, Zerocash supports transparent (like Bitcoin) and shielded transactions, with the latter hiding sender, recipient, and amount. Shielded payments encrypt these values with one-time keys and use zk-SNARK proofs to validate inputs-outputs equality, concealing all on-chain details.
Unspent tokens, introduced in Bitcoin (Nakamoto, 2008), represent transaction outputs that can be used as inputs for future transactions, which in turn create new unspent tokens. As such, they are inherently traceable unless PETs are used (Baldimtsi et al., 2024). Unspent tokens are issued in a decentralized manner, usually as result of the mining process. Their core attributes include an identifier, value (or its commitment), and ownership information. Their recording in a public, transparent ledger prevents double-spending but complicates privacy. Privacy-focused crypto-asset systems employ stealth addresses, ring signatures, confidential transactions, and ZKPs to obscure transaction details and unlink payments (e.g., (Noether et al., 2016; Poelstra, 2016; Miers et al., 2013; Ben Sasson et al., 2014))—these properties are discussed in more detail later. Tokens can also include auxiliary data to support the evaluation of specific policies (e.g., (Xue et al., 2023; Yu et al., 2024)). For example, Xue et al. (Xue et al., 2023) include information to prove in ZK that the total amount of value transferred and the transaction frequency in a time period is limited. For a detailed discussion of auditability in token-based systems, see Section 4.4.
@@
Narula et al. (Narula et al., 2018) propose zkLedger, a third approach tailored to a specific use case: a bank-intermediated CBDC. It uses a tabular ledger where rows represent transactions and columns represent banks. Each cell contains a Pedersen commitment to hide values, while enabling cryptographic verification. Transactions include ZKPs to prove balance, asset ownership, and consistency with audit records. This structure supports rich privacy-preserving auditing: the system includes an auditor who can issue a rich set of auditing queries—using primitives like sums, moving averages, variance, standard deviation, ratios—and receive answers that are provably consistent with the ledger. As downside, this structure does not scale well with many banks and transactions.
@@ 4.2.2. Ledger State Update
Updating the ledger may reveal sensitive information to other parties.
@@
A generic state update transaction for account-based ledger includes the sender and receiver addresses, the exchanged amount, and the sender digital signature. For privacy-enhanced account-based, a naïve state update would reveal the updated accounts. Therefore, privacy-preserving state updates in account-based ledger usually adopt three strategies (e.g., (Kiayias et al., 2022; Wüst et al., 2022; Fauzi et al., 2019; Bünz et al., 2020; Papadoulis et al., 2024)): to obtain confidentiality, they update a commitment of the new account state; to obtain unlinkability, they always update the public key associated with the (recipient) account; and to obtain anonymity, they update a group of accounts at the same time.
@@
Token-based ledgers store the list of valid (unspent) tokens or burnt (spent) tokens. A spending operation must invalidate an existing unspent token, to prevent double-spending. With Chaumian tokens (e.g., (Chaum, 1983; Dold, 2019; Lovejoy et al., 2023)), payment systems include a withdrawal operation, which creates a new valid token that the payer can spend later in time, and a spend operation. The state of the system is represented by the list of spent tokens, which is updated at spending time; unspent tokens are usually not explicitly stored (if not by payers). Privacy-preserving systems based on Chaumian tokens do not usually guarantee transaction confidentiality (e.g., (Chaum, 1983; Hayes, 1990; Sander and Ta-Shma, 1999; Dold, 2019; Goodell et al., 2023; Friolo et al., 2024)), therefore they still maintain the list of spent tokens. Unlike other approaches, Sander and Ta-Shma (Sander and Ta-Shma, 1999) represent valid tokens in a Merkle tree. The bank does not blind sign tokens; instead, valid tokens come with a hash chain proving membership in the Merkle tree. The Merkle root is publicly available to ensure transparency and auditability. Withdrawal coins can be invalidated if needed, addressing certain attack scenarios (e.g., blackmail). Friolo et al. (Friolo et al., 2024) employ a ZKP-based protocol to achieve unlinkability. Transactions are represented as commitments, and when spent, their opening is revealed along with a ZKP demonstrating that it is linked to a valid, unspent token. In contrast, UTT (Tomescu et al., 2022) ensures transaction confidentiality, and it recurs to nullifiers to track spent coins. Nullifiers enable the bank to detect double-spending without knowing transaction specifics. In (Tomescu et al., 2022), a nullifier is a pseudo-random value computed from the coin’s serial number (and a secret). The system checks their validity through ZKPs and, if valid, stores them in a list for future checks.
@@
With unspent tokens, the payment operation simultaneously invalidates previous tokens and generate new ones. The state of the system is represented by the list of unspent tokens [^9] Since blockchains store transactions in an append-only log, they trace the derivation history of each unspent token. This helps implement a decentralize yet scalable system.. Unspent tokens cannot be easily marked as spent without compromising sender anonymity. Most of the approaches, including CryptoNote (Van Saberhagen, 2013), Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), and their extensions (e.g., (Garman et al., 2017; Androulaki and Karame, 2014; Xue et al., 2023)), let the payer spend a token by using a serial number and a ZKP that indicates that the number is a valid opening to some commitment stored in an unspent token. Monero (Noether et al., 2016) (and its extension Traceable Monero (Li et al., 2021)) uses one-time signatures for unspent tokens, which can be used only once to spend the token. At spending time, the ledger registers a salted hash of the public key. MimbleWimble (and PRCash that reuses its transaction format (Wüst et al., 2019)) uses a binding and hiding commitment scheme (Pedersen, 1992), together with Bulletproof range proofs (Bünz et al., 2018), to define the recipient. By using commitments as outputs, coins act as having their own private key, the knowledge of which, along with the knowledge of their value, enable spending the token. This guarantees anonymity and unlinkability.
@@ 4.2.4. Support for Different Denominations
Chaum (Chaum, 1983) considered a simple system where tokens have implicit value. A few extensions consider the possibility to emit tokens with different denominations (e.g., (Dold, 2019; Tomescu et al., 2022)). For example, Dold (Dold, 2019) in GNU Taler use multiple denomination key pairs for blind-signing coins of different monetary values. Conversely, Tomescu et al. (Tomescu et al., 2022) explicitly represent the token value by concealing it within a commitment. As regards unspent token-based ledgers, Zerocoin does not support arbitrary denominations (Miers et al., 2013). This design simplifies the ZKP used for anonymity but introduces usability limitations, as users must split or combine coins when making payments. Zerocash (Ben Sasson et al., 2014) supports arbitrary denominations by using homomorphic commitments, allowing users to transact any amount while preserving privacy. The remaining approaches allow tokens of different denominations, as each unspent token internally holds its value (or a commitment to it). Among them, CryptoNote (Van Saberhagen, 2013) does not hide transaction amounts, making it vulnerable to statistical analysis—e.g., identifying change as the smaller output. Using multiple change addresses can obscure amounts but leads to proliferation of dust transactions (Noether et al., 2016). Hiding values with commitments prevents such analysis.

### v3

@@ 4.1.4. Trust Assumptions
The solutions that introduce trust assumptions do it for different purposes. The approaches using zk-SNARK as ZKP require a trusted setup to initialize and correctly share some cryptographic parameters (e.g., (Miers et al., 2013; Garman et al., 2017; Ben Sasson et al., 2014; Xue et al., 2023)). A few works require a trusted central bank to issue coins that can be later exchanged anonymously (e.g., (Goodell et al., 2023; Friolo et al., 2026)). Elfadul et al. (Elfadul et al., 2024) assume the presence of a regulatory supervision that grants permissions to users to perform transaction; this centralized authority could asymmetrically revoke such authorization. Contributions that explore special forms of auditability related to revocable anonymity (e.g. (Wüst et al., 2019; Lin et al., 2020; Xue et al., 2023; Yu et al., 2024)), trapdoor (e.g., (Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020; Elfadul et al., 2024)), or user tracing (e.g., (Garman et al., 2017)) require trust in the tracing authority that uncovers transaction details only when necessary. PEReDi (Kiayias et al., 2022) reduces trust requirements by relying on quorum of centralized entities that must agree to perform specific operations (such as decrypting transactions).
@@ 4.1.5. Interaction Models
Privacy-preserving digital payment systems differ in whether transactions require direct interaction between payer and payee. Interactive payment protocols need active cooperation during the transaction process. This is common in account-based models where recipients help in constructing or approving the updated account state, via ZKPs or digital signatures. For example, PEReDi (Kiayias et al., 2022) and Platypus (Wüst et al., 2022) enforce synchronized interaction: both parties participate in updating the ledger with proofs and commitments, ensuring mutual accountability and preventing attacks like dusting or replay. These systems may offer enhanced control and auditability but at the cost of real-time coordination and increased communication overhead. Similarly, Chaumian token systems often require immediate token claiming to prevent double spending (e.g., (Dold, 2019)). In contrast, non-interactive payment protocols let the payer unilaterally create and broadcast transactions to the network. Many token-based systems, especially unspent token models, follow this approach (e.g., (Ben Sasson et al., 2014; Van Saberhagen, 2013; Noether et al., 2016; Poelstra, 2016)). This improves user convenience and enables asynchronous payments, although receivers may need periodic ledger scanning or local computation. [^14] Off-chain communication can alleviate the burden of scanning the ledger, but current protocols lack this feature. Some account-based protocols approximate non-interactive behavior via delayed claiming. For instance, Sarencheh et al. (Sarencheh et al., 2025) propose a two stage and non-interactive variant of PEReDi: first, the sender interacts with issuers to update his account (transfer stage); then, the receiver claims funds using enclosed ZKP and credits his account (claim stage). Overall, interactive designs enable rich policy enforcement and conditional privacy, whereas non-interactive models better suit high-latency or asynchronous networks.
@@ 4.2.1. Ledger Data Models
Account-based ledgers associate a balance with each account address, which is debited or credited by transactions. This simplifies state management and enables simple definitions of policies based on cumulative balances. Nonetheless, it requires safeguards against replay and double-spending attacks, often implemented by storing a sequence number (or nonce) in each account and updating it with every transaction. Since the balance of an account is explicitly represented, achieving (pseudo-)anonymity and unlinkability is challenging, especially in systems where users cannot freely generate multiple addresses (e.g., (Guan et al., 2022)). Privacy-preserving account-based solutions replace plain-text balances with commitments, achieving confidentiality of balances and transactions (e.g., (Fauzi et al., 2019; Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022; Sarencheh et al., 2025)) [^15] While in the first generation of systems accounts are kept centralized and private, second and third generation systems are often public, enabling inspection and verification.. Liu et al. (Liu et al., 2022) propose an account-based CBDC with sender and receiver anonymity but with no unlinkability: The system follows a two-tier model where the central bank can identify users via encrypted identifiers, while commercial banks cannot: users use blind signatures by the central bank for account generation and withdrawals and prove validity of their operations using ZKPs. To guarantee anonymity while ensuring unlinkability, most of the approaches change users’ account address at each transaction and use an authenticator to prove correctness. The authenticator can be implemented, e.g., using digital signatures by trusted or centralized authorities (e.g., (Kiayias et al., 2022)) or ZKPs (e.g., (Wüst et al., 2022; Fauzi et al., 2019; Bünz et al., 2020; Papadoulis et al., 2024)). Platypus (Wüst et al., 2022) rely on ZKPs; conversely, in PEReDi (Kiayias et al., 2022), account states must be (blindly) signed by a quorum of maintainers, and since each transaction reveals prior account snapshots to construct new authenticators, randomized signatures are used to prevent linkability. Quisquis (Fauzi et al., 2019) and AQQUA (Papadoulis et al., 2024) use updatable public keys to update accounts in a decentralized manner. The account-based model lends itself well to auditability implemented in third generation systems, as it allows simple definition of policies on holding limits (e.g., (Gross et al., 2021; Wüst et al., 2022; Papadoulis et al., 2024)) or spending and receiving limits (e.g., (Papadoulis et al., 2024)). Gross et al. (Gross et al., 2021) includes, in the account state, an accumulator of all spending in the current epoch, enabling policies on this information. AQQUA (Papadoulis et al., 2024) includes commitments to three different counters (i.e., balance, total amounts sent and received). Conversely, Platypus (Wüst et al., 2022) leaves the specification of the additional information to be application dependent.
@@
Chaumian tokens (Chaum, 1983) use blind signatures and were first deployed in centralized, bank-issued electronic cash schemes (i.e., first generation systems). When issuing a token, the bank blindly signs a random value provided by the payer, allowing the payer to later unblind it for spending (ensuring sender anonymity). These tokens rely on a trusted authority to issue and verify their validity; to this end, the authority must hold an ever increasing list of spent tokens that might introduce scalability challenges. Some implementations introduce expiration to limit circulation (e.g., (Dold, 2019; Lovejoy et al., 2023)). Chaumian-inspired decentralized privacy-preserving systems of the second generation (e.g., Monero (Noether et al., 2016), Zerocoin (Miers et al., 2013), and Zerocash (Ben Sasson et al., 2014)) eliminate the need of a central issuer using a combination of different cryptographic techniques. For example, Zerocash supports transparent (like Bitcoin) and shielded transactions, with the latter hiding sender, recipient, and amount. Shielded payments encrypt these values with one-time keys and use zk-SNARK proofs to validate inputs-outputs equality, concealing all on-chain details.
Unspent tokens, introduced in Bitcoin (Nakamoto, 2008), represent transaction outputs that can be used as inputs for future transactions, which in turn create new unspent tokens. As such, they are inherently traceable unless PETs are used (Baldimtsi et al., 2024). Unspent tokens are issued in a decentralized manner, usually as result of the mining process. Their core attributes include an identifier (e.g., transaction hash and output index), value (or its commitment), and ownership information. Their recording in a public, transparent ledger prevents double-spending but complicates privacy. Privacy-focused crypto-asset systems employ stealth addresses, ring signatures, confidential transactions, and ZKPs to obscure transaction details and unlink payments (e.g., (Noether et al., 2016; Poelstra, 2016; Miers et al., 2013; Ben Sasson et al., 2014))—these properties are discussed in more detail later. Tokens can also include auxiliary data to support the evaluation of specific policies (e.g., (Xue et al., 2023; Yu et al., 2024)) for third generation systems. For example, Xue et al. (Xue et al., 2023) include data to prove in ZK that the total value transferred and the transaction frequency in a time period is limited. Section 4.4 further discusses the auditability of token-based systems.
@@
Narula et al. (Narula et al., 2018) propose zkLedger, a third approach tailored to a specific use case: a bank-intermediated CBDC. It uses a tabular ledger where rows represent transactions and columns represent banks. Each cell contains a Pedersen commitment to hide values, while enabling cryptographic verification. Transactions include ZKPs to prove balance, asset ownership, and consistency with audit records. This structure supports rich privacy-preserving auditing: this third generation system includes an auditor who can issue a rich set of auditing queries—using primitives like sums, moving averages, variance, standard deviation, ratios—and receive answers that are provably consistent with the ledger. As a downside, the tabular structure scales poorly with the number of banks and transactions, since the ledger grows monotonically. MiniLedger (Chatzigiannis and Baldimtsi, 2021) adopts a similar tabular abstraction but addresses scalability by replacing commitments with (twisted ElGamal) encryptions and introducing cryptographically verifiable ledger pruning. MiniLedger allows old transaction rows to be compacted into digests without losing auditability: banks can later selectively disclose individual transactions, aggregates, or balances together with ZKPs, even after pruning. This makes MiniLedger better suited for long-running inter-bank settlement scenarios.
@@ 4.2.2. Ledger State Update
Updating the ledger may reveal sensitive information when it exposes which state elements are created, updated, or consumed to non-involved parties (e.g., globally on public ledgers or to intermediaries), enabling linkability through update patterns (even if values are hidden). Privacy-preserving systems rely on commitments, cryptographic accumulators (e.g., Merkle trees), or tags such as nullifiers, to validate state transitions and prevent double spending while minimizing information leakage.
@@
A generic state update transaction for account-based ledger includes the sender and receiver addresses, the exchanged amount, and the sender digital signature. For privacy-enhanced account-based, a naïve state update would reveal the updated accounts. Therefore, privacy-preserving state updates in account-based ledger usually adopt three strategies (e.g., (Kiayias et al., 2022; Wüst et al., 2022; Fauzi et al., 2019; Bünz et al., 2020; Papadoulis et al., 2024)): to obtain confidentiality, they update a commitment to the new account state; to obtain unlinkability, they always update the public key associated with the (recipient) account; and to obtain anonymity, they update a group of accounts at the same time.
@@
Token-based ledgers store the list of valid (unspent) tokens or burnt (spent) tokens. A spending operation must invalidate an existing unspent token, to prevent double-spending. With Chaumian tokens (e.g., (Chaum, 1983; Dold, 2019; Lovejoy et al., 2023)), payment systems include a withdrawal operation (or mint transaction), which creates a new valid token that the payer can spend later in time, and a spend operation (or pour transaction). The state of the system is represented by the list of spent tokens, which is updated at spending time; unspent tokens are usually not explicitly stored (if not by payers). Privacy-preserving systems based on Chaumian tokens do not usually guarantee transaction confidentiality (e.g., (Chaum, 1983; Hayes, 1990; Sander and Ta-Shma, 1999; Dold, 2019; Goodell et al., 2023; Friolo et al., 2026)), therefore they still maintain the list of spent tokens. Unlike other approaches, Sander and Ta-Shma (Sander and Ta-Shma, 1999) represent valid tokens in a Merkle tree. The bank does not blind sign tokens; instead, valid tokens come with a hash chain proving membership in the Merkle tree. The Merkle root is publicly available to ensure transparency and auditability. Withdrawal coins can be invalidated if needed, addressing certain attack scenarios (e.g., blackmail). To hide amounts, Monero wraps each amount in a Pedersen commitment, and proves via a Bulletproof range ZKP that each value lies in a valid range (no negative or overflow). The additive properties of Pedersen commitments allows proving input-output equality without revealing any value. Friolo et al. (Friolo et al., 2026) employ a ZKP-based protocol to achieve unlinkability: Transactions are represented as commitments, and when spent, their opening is revealed along with a ZKP demonstrating that it is linked to a valid, unspent token. In contrast, UTT (Tomescu et al., 2022) ensures transaction confidentiality, and it recurs to nullifiers to track spent coins. Nullifiers enable the bank to detect double-spending without knowing transaction specifics. In (Tomescu et al., 2022), a nullifier is a pseudo-random value computed from the coin’s serial number (and a secret). The system checks their validity through ZKPs and, if valid, stores them in a list for future checks.
@@
With unspent tokens, the payment operation simultaneously invalidates previous tokens and generate new ones. The state of the system is represented by the list of unspent tokens [^16] Since blockchains store transactions in an append-only log, they trace the derivation history of each unspent token.. Unspent tokens cannot be easily marked as spent without compromising sender anonymity. Most of the approaches, including CryptoNote (Van Saberhagen, 2013), Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), and their extensions (e.g., (Garman et al., 2017; Androulaki and Karame, 2014; Xue et al., 2023)), let the payer spend a token by using a serial number and a ZKP that indicates that the number is a valid opening to some commitment stored in an unspent token. Monero (Noether et al., 2016) (and its extension Traceable Monero (Li et al., 2021)) uses one-time signatures for unspent tokens, which can be used only once to spend the token. At spending time, the ledger registers a salted hash of the public key. MimbleWimble (and PRCash that reuses its transaction format (Wüst et al., 2019)) uses a binding and hiding commitment scheme (Pedersen, 1992), together with Bulletproof range proofs (Bünz et al., 2018), to define the recipient. By using commitments as outputs, coins act as having their own private key, the knowledge of which, along with the knowledge of their value, enable spending the token. This guarantees anonymity and unlinkability.
@@ 4.2.4. Support for Different Denominations
Chaum (Chaum, 1983) considered a simple system where tokens have implicit value. A few extensions consider the possibility to emit tokens with different denominations (e.g., (Dold, 2019; Tomescu et al., 2022)). For example, Dold (Dold, 2019) in GNU Taler uses multiple denomination key pairs for blind-signing coins of different monetary values. Conversely, Tomescu et al. (Tomescu et al., 2022) explicitly represent the token value by concealing it within a commitment. As regards unspent token-based ledgers, Zerocoin does not support arbitrary denominations (Miers et al., 2013). This design simplifies the ZKP used for anonymity but introduces usability limitations, as users must split or combine coins when making payments. Zerocash (Ben Sasson et al., 2014) supports arbitrary denominations by using homomorphic commitments, allowing users to transact any amount while preserving privacy. The remaining approaches allow tokens of different denominations, as each unspent token internally holds its value (or a commitment to it). Among them, CryptoNote (Van Saberhagen, 2013) does not hide transaction amounts, making it vulnerable to statistical analysis—e.g., identifying change as the smaller output. Using multiple change addresses can obscure amounts but leads to proliferation of dust transactions (Noether et al., 2016). Hiding values with commitments prevents such analysis.